Main INDEX
Monthly INDEX
PREV
NEXT
User name R. Michaels
Log entry time 13:05:16 on April 5,2008
Entry number 230053
This entry is a followup to: 229917
keyword=helicity scalers fixed (L-HRS)
I fixed the helicity gating on L-HRS scalers. The 4th input of
the control inputs on the scaler is a veto. It must be the MPS
with correct polarity. I changed this from the NIM-bar output of
a logic module to the NIM output, after verifying on a scope.
Since all this was fine 2 weeks ago, I surmise that someone flipped
the polarity of MPS in the past two weeks, probably to fix the
helicity info in the datastream. My change only affects scalers
and they look fine now.